Calculators were purchased at ​$55 per dozen and sold at ​$15 for three calculators. What is the profit on six dozen​ calculators?
Answer by addingup(3677)

Bought in groups of 12 and sold in groups of 3. How many groups of 3 in 12?
Divide: 12/3 = ? Use your calculator
:
Sold: 12 at 15*4 = 60
Bought: 12 at = 55
Profit for 12(one dozen): 60-55 = 5
Profit for 6 dozen: 6*5 = 30